OpenStudy (anonymous):

how to identify the vertex and the x and y intercept of h(x)=4x^2-4x+21

OpenStudy (unklerhaukus):

the y intercept is when x=0 the x intercept is when y=0 the vertex is when the first derivative is zero

OpenStudy (anonymous):

dont you have to change it into standard form, h(x)=(x-k)+h, first?

OpenStudy (unklerhaukus):

dont do that

OpenStudy (anonymous):

why?

OpenStudy (unklerhaukus):

the equation is already in standard form

OpenStudy (unklerhaukus):

\[h(x)=4x^2-4x+21\] lets find the \(y\)-intercept setting \(x=0\) \[h(0)=4(0)^2-4(0)+21\] \[h(0)=\]

OpenStudy (anonymous):

so it's 21

OpenStudy (anonymous):

but how to find the vertex??

OpenStudy (unklerhaukus):

the \(y\)-intercept is \((0,21)\)

OpenStudy (unklerhaukus):

\[h(x)=4x^2-4x+21\] \[h'(x)=\qquad\qquad\qquad=0\]

OpenStudy (anonymous):

ummm so would i do complete the square?

OpenStudy (unklerhaukus):

take the first derivative

OpenStudy (anonymous):

haha im sorry but what do you mean when you say derivative??

OpenStudy (unklerhaukus):

oh, ok. well have to find the vertex a different way then

OpenStudy (unklerhaukus):

for \[h(x)=ax^2+bx+c\] the x-coordinate is at \[x=\frac{-b}{2a}\]

OpenStudy (anonymous):

i got 1/2

OpenStudy (unklerhaukus):

what is the y coordinate of that point?

OpenStudy (anonymous):

ummm im not sure how to find it...

OpenStudy (unklerhaukus):

you found the x coordinate of the vertex, is 1/2 to find the y coordinate substitute this into the equation h(x)=4x^2-4x+21 h(1/2)=4(1/2)^2+4(1/2)+21

OpenStudy (anonymous):

i got 20

OpenStudy (anonymous):

Thank you very much for the help.